Two particles, A and B, are in uniform circular motion about a common center. The acceleration of particle A is 4.7 times that of particle B. Particle B takes 2.4 times as long for a rotation as particle A. The ratio of the radius of the motion of particle A to that of particle B is closest to:

A) 2.0

B) 3.8

C)0.82

D)0.51

E)11
Answer #1

aA = 4.7 aB

TB = 2.4 TA

a = mrw^2 = mr(2pi/T)^2

aA/aB = (rA/rB)(TB/TA)^2

4.7 = (rA/rB) (2.4)^2

rA/rB = 0.82

Correct option is (C)